OpenStudy (anonymous):

For a ground temperature of 15 degree Celsius, the air temperature (t) at an altitude of 500m (h) is 10 degree Celsius. At 2000 m, the air temperature is -5 degree Celsius. Use the ordered pairs (h,t). a. Find a linear equation for these data points. b. Use this linear equation to find the air temperature at 1500m.

OpenStudy (anonymous):

y=mx+b

OpenStudy (mertsj):

You have been given two ordered pairs: (500,10) and (2000,-5) Can you find the slope of this line?

OpenStudy (anonymous):

No I also need help with that.

OpenStudy (mertsj):

Subtract the second numbers of the ordered pairs: 10-(-5)=15 Subtract the first numbers of the ordered pairs: (500-2000)=-1500 Put the first number over the second: 15/-1500 =-1/100 That is the slope.

OpenStudy (mertsj):

Now substitute the slope and one of the ordered pairs into y = mx + b and find b. Then you will have the equation of your line.

OpenStudy (mertsj):

Can you do that?

OpenStudy (anonymous):

im doing it now hang on please

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

What did you get for b?

OpenStudy (anonymous):

wait, I think I did something wrong lemmy double check.

OpenStudy (mertsj):

double check granted.

OpenStudy (anonymous):

can you pick either one (15/-1500) or (-1/100) as the x and the y?

OpenStudy (mertsj):

-1/100 is m, the slope. For x and y you may use either ordered pair you choose.

OpenStudy (anonymous):

I got -15 = (1500)(-1/100) + b

OpenStudy (mertsj):

\[10=\frac{-1}{100}(500)+b\] \[10=-5+b\] \[15=b\]

OpenStudy (mertsj):

Solve for b

OpenStudy (anonymous):

oh I see now. Thank you for being patient with me.

OpenStudy (mertsj):

\[y=\frac{-1}{100}x+15\]

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

Now replace x with 1500 to find the temperature at 1500 m